# Pricing Options on Forwards in Function-Valued Affine Stochastic Volatility Models

### v1

@@ 4.3. Pricing in a pure-jump Lévy driven stochastic volatility model
Let $\mathcal{A}$ be the generator of the left-shift semigroup and $(\widetilde{W}_{t})_{t\geq 0}$ a cylindrical Brownian motion under $\mathbb{Q}$ as in Section 3. Consider a compound Poisson process $(J_{t})_{t\geq 0}$ with deterministic jump directions $D$ and intensity $\beta>0$:
$J_{t}=\sum_{i=1}^{N_{t}}D,\qquad N_{t}\sim\mathrm{Poisson}(\beta t).$
@@
$\mathop{}\!\textnormal{d}X_{t}$ $=\big{(}\mathcal{A}X_{t}+D^{1/2}\mathsf{Y}_{t}D^{1/2}\Upsilon^{\vartheta}\big{)}\,\mathop{}\!\textnormal{d}t+D^{1/2}\mathsf{Y}_{t}^{1/2}\,\mathop{}\!\textnormal{d}\widetilde{W}_{t},\qquad X_{0}=x\in H_{w},$
$\mathop{}\!\textnormal{d}\mathsf{Y}_{t}$ $=\beta D\,\mathop{}\!\textnormal{d}t+\mathop{}\!\textnormal{d}J_{t},\qquad\mathsf{Y}_{0}=y\in\mathcal{H}_{w}^{+},$ (4.4)
with $\Upsilon^{\vartheta}=-\tfrac{1}{2}S^{*}(\vartheta)h_{0}$ as in (3.7).
@@
$\partial_{t}\phi(t,u)=F(\psi_{2}(t,u)),\qquad\psi_{1}(t,u)=u_{1}+\mathcal{A}^{*}\int_{0}^{t}\psi_{1}(s,u)\,\mathop{}\!\textnormal{d}s,\qquad\partial_{t}\psi_{2}(t,u)=R(\psi_{1}(t,u),\psi_{2}(t,u)),$
@@
$R(h,u)=-\tfrac{1}{2}\,(D^{1/2}h)^{\otimes 2}-D^{1/2}h\otimes D^{1/2}\Upsilon^{\vartheta},\qquad F(u)=\beta\big{(}\mathop{}\!\textnormal{e}^{-\langle D,u\rangle}-1\big{)},$
@@
Let $w(x)=e^{\alpha x}$ with $\alpha>0$ and $(f_{n})_{n\in\mathbb{N}}$ as above. For $u_{1}=(\nu+i\lambda)\,u_{\vartheta}$ with $\nu>1$ and $\lambda\in\mathbb{R}$, the Riccati solution satisfies
$\psi_{1}(t,u)(x)$ $=(\nu+i\lambda)\!\left(1+\frac{1}{\alpha}\big{(}1-e^{-\alpha(x\wedge t)}\big{)}+\frac{1}{\alpha}e^{-\alpha t}\,\mathbf{1}_{\{x\geq t\}}\big{(}1-e^{-\alpha(\vartheta\wedge x-t)}\big{)}\right),$ (4.5)
$\psi_{2}(t,u)$ $=-\frac{(\nu+i\lambda)^{2}}{4}\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\frac{1}{ij}\,\langle f_{i},\cdot\rangle_{w}\,f_{j}\int_{0}^{t}f_{i}(s+\vartheta)f_{j}(s+\vartheta)\,\mathop{}\!\textnormal{d}s$
$\quad\;-\frac{(\nu+i\lambda)}{2}\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\frac{1}{ij}\,c_{j}(\vartheta)\,\langle f_{i},\cdot\rangle_{w}\,f_{j}\int_{0}^{t}f_{i}(s+\vartheta)\,\mathop{}\!\textnormal{d}s,$ (4.6)
where $D^{1/2}=\tfrac{1}{\sqrt{2}}\sum_{n\geq 1}\frac{1}{n}\,f_{n}\otimes f_{n}$ and $c_{j}(\vartheta):=\langle f_{j},S^{*}(\vartheta)h_{0}\rangle_{w}$, i.e., $D^{1/2}\Upsilon^{\vartheta}=\tfrac{1}{\sqrt{2}}\sum_{j\geq 1}\frac{1}{j}\,c_{j}(\vartheta)\,f_{j}$. Moreover,
$\langle D,\psi_{2}(t,u)\rangle$ $=-\frac{(\nu+i\lambda)^{2}}{8}\sum_{n=1}^{\infty}\frac{1}{n^{4}}\int_{0}^{t}f_{n}(s+\vartheta)^{2}\,\mathop{}\!\textnormal{d}s\;-\;\frac{(\nu+i\lambda)}{4}\sum_{n=1}^{\infty}\frac{c_{n}(\vartheta)}{n^{4}}\int_{0}^{t}f_{n}(s+\vartheta)\,\mathop{}\!\textnormal{d}s,$ (4.7)
@@
$\phi(t,u)$ $=\beta t-\beta\int_{0}^{t}\exp{\left\{\frac{1}{8}(\nu+\mathrm{i}\lambda)^{2}\sum_{k=1}^{\infty}\frac{1}{k^{4}}\int_{0}^{s}f_{k}(l+\vartheta)\left(f_{k}(l+\vartheta)+\frac{2c_{k}(\vartheta)}{\nu+\mathrm{i}\lambda}\right)\,\mathop{}\!\textnormal{d}l\right\}}\,\mathop{}\!\textnormal{d}s\,.$ (4.8)
@@
$\partial_{t}\psi_{2}=-\tfrac{1}{2}(D^{1/2}\psi_{1})^{\otimes 2}-D^{1/2}\psi_{1}\otimes D^{1/2}\Upsilon^{\vartheta}$ yields (4.6). For $\langle D,\psi_{2}\rangle=\sum_{n\geq 1}\langle Df_{n},\psi_{2}f_{n}\rangle_{w}$, use $Df_{n}=\tfrac{1}{2n^{2}}f_{n}$ and orthonormality to obtain (4.7). Finally, $\partial_{t}\phi=F(\psi_{2})$ with $F(u)=-\beta\big{(}e^{-\langle D,u\rangle}-1\big{)}$ gives (4.8). ∎
@@
$\mathbb{E}_{\mathbb{Q}}\big{[}(F(t,t+\vartheta)-K)^{+}\mid\mathcal{F}_{0}\big{]}$ $=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\Big{(}-\phi(t,u)+\langle X_{0},\psi_{1}(t,u)\rangle_{w}-\langle Y_{0},\psi_{2}(t,u)\rangle\Big{)}\,\mathop{}\!\textnormal{d}\lambda$ (4.9)
$=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\Big{(}-\beta t+\beta\int_{0}^{t}\mathop{}\!\textnormal{e}^{-\langle D,\psi_{2}(s,u)\rangle}\,\mathop{}\!\textnormal{d}s$
$\qquad\qquad+(\nu+i\lambda)\,X_{0}(t+\vartheta)-\langle Y_{0},\psi_{2}(t,u)\rangle\Big{)}\,\mathop{}\!\textnormal{d}\lambda,$
@@
$\langle D,\psi_{2}(s,u)\rangle=-\frac{(\nu+i\lambda)^{2}}{8}\sum_{n=1}^{\infty}\frac{1}{n^{4}}\int_{0}^{s}f_{n}(r+\vartheta)^{2}\,\mathop{}\!\textnormal{d}r-\frac{(\nu+i\lambda)}{4}\sum_{n=1}^{\infty}\frac{c_{n}(\vartheta)}{n^{4}}\int_{0}^{s}f_{n}(r+\vartheta)\,\mathop{}\!\textnormal{d}r,$
and $c_{n}(\vartheta):=\langle f_{n},S^{*}(\vartheta)h_{0}\rangle_{w}$.
@@
$\mathbb{E}_{\mathbb{Q}}\!\big{[}(F(t,t+\vartheta)-K)^{+}\mid\mathcal{F}_{0}\big{]}$ $\approx\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\Big{(}-\beta t+\beta\int_{0}^{t}\mathop{}\!\textnormal{e}^{-\langle D,\psi_{2}^{(N)}(s,u)\rangle}\,\mathop{}\!\textnormal{d}s$ (4.10)
$\qquad\qquad\qquad\qquad+(\nu+i\lambda)X_{0}(t+\vartheta)-\langle Y_{0},\psi_{2}^{(N)}(t,u)\rangle\Big{)}\,\mathop{}\!\textnormal{d}\lambda,$
@@
$\langle D,\psi_{2}^{(N)}(s,u)\rangle=-\frac{(\nu+i\lambda)^{2}}{8}\sum_{n=1}^{N}\frac{1}{n^{4}}\int_{0}^{s}f_{n}(r+\vartheta)^{2}\,\mathop{}\!\textnormal{d}r-\frac{(\nu+i\lambda)}{4}\sum_{n=1}^{N}\frac{c_{n}(\vartheta)}{n^{4}}\int_{0}^{s}f_{n}(r+\vartheta)\,\mathop{}\!\textnormal{d}r.$
The function $f_{n}$ can be efficiently evaluated by iteration, see Appendix A, and consequently the integrals $\int_{0}^{s}f_{n}(r+\vartheta)\,\mathop{}\!\textnormal{d}r$ and $\int_{0}^{s}f_{n}(r+\vartheta)^{2}\,\mathop{}\!\textnormal{d}r$ can be calculated by using Gaussian quadrature method.
@@
Figure 2. Monte Carlo v.s. analytical formula, with $Y_{0}=D,h_{0}=1,\alpha=0.1,T_{0}=1,N=5$, strike $K=1$ and the intensity $\beta=1$.
@@ 4.4. Pricing in the operator-valued Barndorff–Nielsen–Shephard stochastic volatility model
Let $\mathcal{A}$, $D$, $(\widetilde{W}_{t})_{t\geq 0}$, and $(J_{t})_{t\geq 0}$ be as in Example 4.3, with $J_{t}=\sum_{i=1}^{N_{t}}D$ a compound Poisson process of intensity $\beta>0$, independent of $\widetilde{W}$. Let $\mathcal{H}_{w}$ be the space of self-adjoint Hilbert–Schmidt operators on $H_{w}$ and $\mathcal{H}_{w}^{+}$ its positive cone. We consider the operator-valued BNS dynamics
$\mathop{}\!\textnormal{d}X_{t}$ $=\big{(}\mathcal{A}X_{t}+D^{1/2}\mathsf{Y}_{t}D^{1/2}\Upsilon^{\vartheta}\big{)}\,\mathop{}\!\textnormal{d}t+D^{1/2}\mathsf{Y}_{t}^{1/2}\,\mathop{}\!\textnormal{d}\widetilde{W}_{t},\qquad X_{0}=x\in H_{w},$
$\mathop{}\!\textnormal{d}\mathsf{Y}_{t}$ $=\big{(}Q\mathsf{Y}_{t}+\mathsf{Y}_{t}Q^{*}\big{)}\,\mathop{}\!\textnormal{d}t+\mathop{}\!\textnormal{d}J_{t},\qquad\mathsf{Y}_{0}=y\in\mathcal{H}_{w}^{+},$ (4.11)
@@
$\partial_{t}\phi(t,u)=F(\psi_{2}(t,u)),\qquad\psi_{1}(t,u)=u_{1}+\mathcal{A}^{*}\int_{0}^{t}\psi_{1}(s,u)\,\mathop{}\!\textnormal{d}s,\qquad\partial_{t}\psi_{2}(t,u)=R(\psi_{1}(t,u),\psi_{2}(t,u)),$
@@
$R(h,u)=uQ^{*}+Qu-\tfrac{1}{2}\,(D^{1/2}h)^{\otimes 2}-D^{1/2}h\otimes D^{1/2}\Upsilon^{\vartheta},\qquad F(u)=\beta\big{(}\mathop{}\!\textnormal{e}^{-\langle D,u\rangle}-1\big{)},$
@@
Let $w(x)=e^{\alpha x}$ with $\alpha>0$ and $(f_{n})_{n\in\mathbb{N}}$ be the orthonormal basis from (4.2). Assume the OU drift is diagonal,
@@
and set $D^{1/2}=\tfrac{1}{\sqrt{2}}\sum_{n\geq 1}\tfrac{1}{n}\,f_{n}\otimes f_{n}$, $c_{j}(\vartheta):=\langle f_{j},S^{*}(\vartheta)h_{0}\rangle_{w}$ (so $D^{1/2}\Upsilon^{\vartheta}=\tfrac{1}{\sqrt{2}}\sum_{j\geq 1}\tfrac{1}{j}c_{j}(\vartheta)\,f_{j}$). For $u_{1}=(\nu+i\lambda)u_{\vartheta}$ with $\nu>1$, $\lambda\in\mathbb{R}$, the complex Riccati solution to (4.4) satisfies
$\psi_{1}(t,u)(x)$ $=(\nu+i\lambda)\!\left(1+\frac{1}{\alpha}\big{(}1-e^{-\alpha(x\wedge t)}\big{)}+\frac{1}{\alpha}e^{-\alpha t}\,\mathbf{1}_{\{x\geq t\}}\big{(}1-\mathop{}\!\textnormal{e}^{-\alpha(\vartheta\wedge x-t)}\big{)}\right),$ (4.12)
$\psi_{2}(t,u)f_{n}$ $=-\sum_{j\geq 1}\int_{0}^{t}\mathop{}\!\textnormal{e}^{-(a_{n}+a_{j})(t-s)}\Big{[}\frac{(\nu+i\lambda)^{2}}{4\,nj}\,f_{n}(s+\vartheta)f_{j}(s+\vartheta)+\frac{(\nu+i\lambda)}{2\,nj}\,c_{j}(\vartheta)\,f_{n}(s+\vartheta)\Big{]}\mathop{}\!\textnormal{d}sf_{j},$ (4.13)
@@
$\phi(t,u)$ $=\beta t-\beta\int_{0}^{t}\exp\big{(}\frac{1}{8}(\nu+\mathrm{i}\lambda)^{2}\sum_{n=1}^{\infty}\frac{1}{n^{4}}e^{\frac{s}{n^{2}}}\int_{0}^{s}\mathop{}\!\textnormal{e}^{-\frac{l}{n^{2}}}f_{n}(l+\vartheta)\left(f_{n}(l+\vartheta)+\frac{2c_{n}(\vartheta)}{\nu+\mathrm{i}\lambda}\right)\,\mathop{}\!\textnormal{d}l\big{)}\,\mathop{}\!\textnormal{d}s.$ (4.14)
@@
$\langle D,\psi_{2}(t,u)\rangle=-\int_{0}^{t}\Big{[}\frac{(\nu+i\lambda)^{2}}{8}\sum_{n\geq 1}\frac{1}{n^{4}}f_{n}(s+\vartheta)^{2}+\frac{(\nu+i\lambda)}{4}\sum_{n\geq 1}\frac{c_{n}(\vartheta)}{n^{4}}f_{n}(s+\vartheta)\Big{]}\mathop{}\!\textnormal{e}^{-2a_{n}(t-s)}\,\mathop{}\!\textnormal{d}s.$ (4.15)
@@
Since $\psi_{1}$ solves $\psi_{1}(t,u)=u_{1}+\mathcal{A}^{*}\int_{0}^{t}\psi_{1}(s,u)\,\mathop{}\!\textnormal{d}s$ with $u_{1}=(\nu+i\lambda)u_{\vartheta}$, we have the mild form $\psi_{1}(t,u)=S^{*}(t)u_{1}$, which yields (4.12) exactly as in Lemma 4.1.
@@
$\psi_{2}(t)=\int_{0}^{t}\mathop{}\!\textnormal{e}^{(t-s)\mathcal{L}_{Q}}\left[\,-\frac{1}{2}(D^{1/2}\psi_{1}(s))^{\otimes 2}-D^{1/2}\psi_{1}(s)\otimes D^{1/2}\Upsilon^{\vartheta}\right]\mathop{}\!\textnormal{d}s.$ (4.16)

### v2

@@ 4.3. Pricing in a pure-jump Lévy driven stochastic volatility model
Let $\mathcal{A}$ be the generator of the left-shift semigroup and $(\widetilde{W}_{t})_{t\geq 0}$ a cylindrical Brownian motion under $\mathbb{Q}$ as in Section 3. Consider the compensated compound Poisson martingale $(J_{t})_{t\geq 0}$ with deterministic jump directions $D$ and intensity $\beta>0$:
$J_{t}=(N_{t}-\beta t)D,\qquad N_{t}\sim\mathrm{Poisson}(\beta t).$
@@
$\mathop{}\!\mathrm{d}X_{t}$ $=\big(\mathcal{A}X_{t}+D^{1/2}\mathsf{Y}_{t}D^{1/2}\Upsilon^{\vartheta}\big)\,\mathop{}\!\mathrm{d}t+D^{1/2}\mathsf{Y}_{t}^{1/2}\,\mathop{}\!\mathrm{d}\widetilde{W}_{t},\qquad X_{0}=x\in H_{w},$
$\mathop{}\!\mathrm{d}\mathsf{Y}_{t}$ $=\beta D\,\mathop{}\!\mathrm{d}t+\mathop{}\!\mathrm{d}J_{t},\qquad\mathsf{Y}_{0}=y\in\mathcal{H}_{w}^{+},$ (4.4)
with $\Upsilon^{\vartheta}=-\tfrac{1}{2}S^{*}(\vartheta)h_{0}$ as in (3.7). The numerical experiments in Sections 4.3 and 4.4 specialize to the Lévy-driven case $\mu=0$ and to a compound-Poisson OU/BNS specification with jumps in the fixed direction $D$. Section 4.5 extends the validation to the fully state-dependent jump kernel from Section 2.1.
@@
$\partial_{t}\phi(t,u)=F(\psi_{2}(t,u)),\qquad\psi_{1}(t,u)=u_{1}+\mathcal{A}^{*}\int_{0}^{t}\psi_{1}(s,u)\,\mathop{}\!\mathrm{d}s,\qquad\partial_{t}\psi_{2}(t,u)=R(\psi_{1}(t,u),\psi_{2}(t,u)),$
@@
$R(h,u)=-\tfrac{1}{2}\,(D^{1/2}h)^{\otimes 2}-D^{1/2}h\otimes D^{1/2}\Upsilon^{\vartheta},\qquad F(u)=-\beta\big(\mathop{}\!\mathrm{e}^{-\langle D,u\rangle}-1\big).$
@@
Let $w(x)=\mathop{}\!\mathrm{e}^{\alpha x}$ with $\alpha>0$ and $(f_{n})_{n\in\mathbb{N}}$ as above. For $u_{1}=(\nu+i\lambda)\,u_{\vartheta}$ with $\nu>1$ and $\lambda\in\mathbb{R}$, the Riccati solution satisfies
$\psi_{1}(t,u)(x)$ $=(\nu+i\lambda)\!\left(1+\frac{1}{\alpha}\big(1-\mathop{}\!\mathrm{e}^{-\alpha(x\wedge t)}\big)+\frac{1}{\alpha}\mathop{}\!\mathrm{e}^{-\alpha t}\,\mathbf{1}_{\{x\geq t\}}\big(1-\mathop{}\!\mathrm{e}^{-\alpha(\vartheta\wedge x-t)}\big)\right),$ (4.5)
$\psi_{2}(t,u)$ $=-\frac{(\nu+i\lambda)^{2}}{4}\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\frac{1}{ij}\,\langle f_{i},\cdot\rangle_{w}\,f_{j}\int_{0}^{t}f_{i}(s+\vartheta)f_{j}(s+\vartheta)\,\mathop{}\!\mathrm{d}s$
$\quad\;-\frac{(\nu+i\lambda)}{2}\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\frac{1}{ij}\,c_{j}(\vartheta)\,\langle f_{i},\cdot\rangle_{w}\,f_{j}\int_{0}^{t}f_{i}(s+\vartheta)\,\mathop{}\!\mathrm{d}s,$ (4.6)
where $D^{1/2}=\tfrac{1}{\sqrt{2}}\sum_{n\geq 1}\frac{1}{n}\,f_{n}\otimes f_{n}$ and $c_{j}(\vartheta):=-\tfrac{1}{2}\langle f_{j},S^{*}(\vartheta)h_{0}\rangle_{w}$, i.e., $D^{1/2}\Upsilon^{\vartheta}=\tfrac{1}{\sqrt{2}}\sum_{j\geq 1}\frac{1}{j}\,c_{j}(\vartheta)\,f_{j}$. Moreover,
$\langle D,\psi_{2}(t,u)\rangle$ $=-\frac{(\nu+i\lambda)^{2}}{8}\sum_{n=1}^{\infty}\frac{1}{n^{4}}\int_{0}^{t}f_{n}(s+\vartheta)^{2}\,\mathop{}\!\mathrm{d}s\;-\;\frac{(\nu+i\lambda)}{4}\sum_{n=1}^{\infty}\frac{c_{n}(\vartheta)}{n^{4}}\int_{0}^{t}f_{n}(s+\vartheta)\,\mathop{}\!\mathrm{d}s,$ (4.7)
@@
$\phi(t,u)$ $=\beta t-\beta\int_{0}^{t}\exp\left\{\frac{1}{8}(\nu+\mathrm{i}\lambda)^{2}\sum_{k=1}^{\infty}\frac{1}{k^{4}}\int_{0}^{s}f_{k}(l+\vartheta)\left(f_{k}(l+\vartheta)+\frac{2c_{k}(\vartheta)}{\nu+\mathrm{i}\lambda}\right)\,\mathop{}\!\mathrm{d}l\right\}\,\mathop{}\!\mathrm{d}s\,.$ (4.8)
@@
$\partial_{t}\psi_{2}=-\tfrac{1}{2}(D^{1/2}\psi_{1})^{\otimes 2}-D^{1/2}\psi_{1}\otimes D^{1/2}\Upsilon^{\vartheta}$ yields (4.6). For $\langle D,\psi_{2}\rangle=\sum_{n\geq 1}\langle Df_{n},\psi_{2}f_{n}\rangle_{w}$, use $Df_{n}=\tfrac{1}{2n^{2}}f_{n}$ and orthonormality to obtain (4.7). Finally, $\partial_{t}\phi=F(\psi_{2})$ with $F(u)=-\beta\big(\mathop{}\!\mathrm{e}^{-\langle D,u\rangle}-1\big)$ gives (4.8). ∎
@@
$\mathbb{E}_{\mathbb{Q}}\big[(F(t,t+\vartheta)-K)^{+}\mid\mathcal{F}_{0}\big]$ $=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\Big(-\phi(t,u)+\langle X_{0},\psi_{1}(t,u)\rangle_{w}-\langle Y_{0},\psi_{2}(t,u)\rangle\Big)\,\mathop{}\!\mathrm{d}\lambda$ (4.9)
$=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\Big(-\beta t+\beta\int_{0}^{t}\mathop{}\!\mathrm{e}^{-\langle D,\psi_{2}(s,u)\rangle}\,\mathop{}\!\mathrm{d}s$
$\qquad\qquad+(\nu+i\lambda)\,X_{0}(t+\vartheta)-\langle Y_{0},\psi_{2}(t,u)\rangle\Big)\,\mathop{}\!\mathrm{d}\lambda,$
@@
$\langle D,\psi_{2}(s,u)\rangle=-\frac{(\nu+i\lambda)^{2}}{8}\sum_{n=1}^{\infty}\frac{1}{n^{4}}\int_{0}^{s}f_{n}(r+\vartheta)^{2}\,\mathop{}\!\mathrm{d}r-\frac{(\nu+i\lambda)}{4}\sum_{n=1}^{\infty}\frac{c_{n}(\vartheta)}{n^{4}}\int_{0}^{s}f_{n}(r+\vartheta)\,\mathop{}\!\mathrm{d}r,$
and $c_{n}(\vartheta):=-\tfrac{1}{2}\langle f_{n},S^{*}(\vartheta)h_{0}\rangle_{w}$.
@@
$\mathbb{E}_{\mathbb{Q}}\!\big[(F(t,t+\vartheta)-K)^{+}\mid\mathcal{F}_{0}\big]$ $\approx\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\Big(-\beta t+\beta\int_{0}^{t}\mathop{}\!\mathrm{e}^{-\langle D,\psi_{2}^{(N)}(s,u)\rangle}\,\mathop{}\!\mathrm{d}s$ (4.10)
$\qquad\qquad\qquad\qquad+(\nu+i\lambda)X_{0}(t+\vartheta)-\langle Y_{0},\psi_{2}^{(N)}(t,u)\rangle\Big)\,\mathop{}\!\mathrm{d}\lambda,$
@@
$\langle D,\psi_{2}^{(N)}(s,u)\rangle=-\frac{(\nu+i\lambda)^{2}}{8}\sum_{n=1}^{N}\frac{1}{n^{4}}\int_{0}^{s}f_{n}(r+\vartheta)^{2}\,\mathop{}\!\mathrm{d}r-\frac{(\nu+i\lambda)}{4}\sum_{n=1}^{N}\frac{c_{n}(\vartheta)}{n^{4}}\int_{0}^{s}f_{n}(r+\vartheta)\,\mathop{}\!\mathrm{d}r.$
The function $f_{n}$ can be efficiently evaluated by iteration, see Appendix A. The integrals $\int_{0}^{s}f_{n}(r+\vartheta)\,\mathop{}\!\mathrm{d}r$ and $\int_{0}^{s}f_{n}(r+\vartheta)^{2}\,\mathop{}\!\mathrm{d}r$ are computed numerically via Gauss–Legendre quadrature; in our implementation we use the scipy.integrate.quad routine, which provides adaptive Gaussian quadrature and is sufficient for the smooth integrands arising from the basis functions $f_{n}$.
@@
Figure 3. Conditional Monte Carlo vs. affine formula, with $Y_{0}=D,h_{0}=1,\alpha=0.1,T_{0}=1,N=5$, strike $K=1$ and the intensity $\beta=1$.
@@ 4.4. Pricing in the operator-valued Barndorff–Nielsen–Shephard stochastic volatility model
Let $\mathcal{A}$, $D$, $(\widetilde{W}_{t})_{t\geq 0}$, and $(J_{t})_{t\geq 0}$ be as in Example 4.3, with $J_{t}=(N_{t}-\beta t)D$ the compensated compound Poisson martingale of intensity $\beta>0$, independent of $\widetilde{W}$. Let $\mathcal{H}_{w}$ be the space of self-adjoint Hilbert–Schmidt operators on $H_{w}$ and $\mathcal{H}_{w}^{+}$ its positive cone. We consider the operator-valued BNS dynamics
$\mathop{}\!\mathrm{d}X_{t}$ $=\big(\mathcal{A}X_{t}+D^{1/2}\mathsf{Y}_{t}D^{1/2}\Upsilon^{\vartheta}\big)\,\mathop{}\!\mathrm{d}t+D^{1/2}\mathsf{Y}_{t}^{1/2}\,\mathop{}\!\mathrm{d}\widetilde{W}_{t},\qquad X_{0}=x\in H_{w},$
$\mathop{}\!\mathrm{d}\mathsf{Y}_{t}$ $=\big(Q\mathsf{Y}_{t}+\mathsf{Y}_{t}Q^{*}\big)\,\mathop{}\!\mathrm{d}t+\mathop{}\!\mathrm{d}J_{t},\qquad\mathsf{Y}_{0}=y\in\mathcal{H}_{w}^{+},$ (4.11)
@@
$\partial_{t}\phi(t,u)=F(\psi_{2}(t,u)),\qquad\psi_{1}(t,u)=u_{1}+\mathcal{A}^{*}\int_{0}^{t}\psi_{1}(s,u)\,\mathop{}\!\mathrm{d}s,\qquad\partial_{t}\psi_{2}(t,u)=R(\psi_{1}(t,u),\psi_{2}(t,u)),$
@@
$R(h,u)=uQ^{*}+Qu-\tfrac{1}{2}\,(D^{1/2}h)^{\otimes 2}-D^{1/2}h\otimes D^{1/2}\Upsilon^{\vartheta},\qquad F(u)=-\beta\big(\mathop{}\!\mathrm{e}^{-\langle D,u\rangle}-1\big),$
@@
Let $w(x)=\mathop{}\!\mathrm{e}^{\alpha x}$ with $\alpha>0$ and $(f_{n})_{n\in\mathbb{N}}$ be the orthonormal basis from (4.2). Assume the OU drift is diagonal,
@@
and set $D^{1/2}=\tfrac{1}{\sqrt{2}}\sum_{n\geq 1}\tfrac{1}{n}\,f_{n}\otimes f_{n}$, $c_{j}(\vartheta):=-\tfrac{1}{2}\langle f_{j},S^{*}(\vartheta)h_{0}\rangle_{w}$ (so $D^{1/2}\Upsilon^{\vartheta}=\tfrac{1}{\sqrt{2}}\sum_{j\geq 1}\tfrac{1}{j}c_{j}(\vartheta)\,f_{j}$). For $u_{1}=(\nu+i\lambda)u_{\vartheta}$ with $\nu>1$, $\lambda\in\mathbb{R}$, the complex Riccati solution to (4.4) satisfies
$\psi_{1}(t,u)(x)$ $=(\nu+i\lambda)\!\left(1+\frac{1}{\alpha}\big(1-\mathop{}\!\mathrm{e}^{-\alpha(x\wedge t)}\big)+\frac{1}{\alpha}\mathop{}\!\mathrm{e}^{-\alpha t}\,\mathbf{1}_{\{x\geq t\}}\big(1-\mathop{}\!\mathrm{e}^{-\alpha(\vartheta\wedge x-t)}\big)\right),$ (4.12)
$\psi_{2}(t,u)f_{n}$ $=-\sum_{j\geq 1}\int_{0}^{t}\mathop{}\!\mathrm{e}^{-(a_{n}+a_{j})(t-s)}\Big[\frac{(\nu+i\lambda)^{2}}{4\,nj}\,f_{n}(s+\vartheta)f_{j}(s+\vartheta)+\frac{(\nu+i\lambda)}{2\,nj}\,c_{j}(\vartheta)\,f_{n}(s+\vartheta)\Big]\mathop{}\!\mathrm{d}sf_{j},$ (4.13)
@@
$\phi(t,u)$ $=\beta t-\beta\int_{0}^{t}\exp\big(\frac{1}{8}(\nu+\mathrm{i}\lambda)^{2}\sum_{n=1}^{\infty}\frac{1}{n^{4}}\mathop{}\!\mathrm{e}^{\frac{s}{n^{2}}}\int_{0}^{s}\mathop{}\!\mathrm{e}^{-\frac{l}{n^{2}}}f_{n}(l+\vartheta)\left(f_{n}(l+\vartheta)+\frac{2c_{n}(\vartheta)}{\nu+\mathrm{i}\lambda}\right)\,\mathop{}\!\mathrm{d}l\big)\,\mathop{}\!\mathrm{d}s.$ (4.14)
@@
$\langle D,\psi_{2}(t,u)\rangle=-\int_{0}^{t}\Big[\frac{(\nu+i\lambda)^{2}}{8}\sum_{n\geq 1}\frac{1}{n^{4}}f_{n}(s+\vartheta)^{2}+\frac{(\nu+i\lambda)}{4}\sum_{n\geq 1}\frac{c_{n}(\vartheta)}{n^{4}}f_{n}(s+\vartheta)\Big]\mathop{}\!\mathrm{e}^{-2a_{n}(t-s)}\,\mathop{}\!\mathrm{d}s.$ (4.15)
@@
Since $\psi_{1}$ solves $\psi_{1}(t,u)=u_{1}+\mathcal{A}^{*}\int_{0}^{t}\psi_{1}(s,u)\,\mathop{}\!\mathrm{d}s$ with $u_{1}=(\nu+i\lambda)u_{\vartheta}$, we have the mild form $\psi_{1}(t,u)=S^{*}(t)u_{1}$, which yields (4.12) exactly as in Lemma 4.1.
@@
$\psi_{2}(t)=\int_{0}^{t}\mathop{}\!\mathrm{e}^{(t-s)\mathcal{L}_{Q}}\left[\,-\frac{1}{2}(D^{1/2}\psi_{1}(s))^{\otimes 2}-D^{1/2}\psi_{1}(s)\otimes D^{1/2}\Upsilon^{\vartheta}\right]\mathop{}\!\mathrm{d}s.$ (4.16)